# No License

### v3

@@ 6 Conclusion
In summary, the paper provides a novel perspective for creating a powerful GNN across all levels when the space of node features is uncountable. The central idea is to use pseudometric distances to create soft-injective functions such that distinct inputs may produce similar outputs if and only if the distance between inputs is sufficiently small on some representation. From the results, the SIR-GCN is proposed as the first MPNN instance to emphasize contextualized message transformation, setting it apart from other GNNs. Furthermore, the model is shown to generalize classical GNN methodologies. Despite its simple design, empirical results underscore SIR-GCN as the best-performing MPNN instance that effectively balances model complexity and expressivity. The paper thus contributes to GNN literature by theoretically and empirically demonstrating the necessity of both anisotropic and dynamic messages to enhance GNN expressivity. Future works may consider incorporating SIR-GCN into complex frameworks, such as grouped reversible residual connections (Li et al., 2021), adaptive graph diffusion (Sun et al., 2020), graph stochastic attention (Miao et al., 2022), and hierarchical message-passing (Bodnar et al., 2021), to address the limitations inherent in the MPNN framework and to develop more expressive GNN architectures.
@@ Acknowledgments
The work is supported by the Japan Society for the Promotion of Science through the Grants-in-Aid for Scientific Research Program (KAKENHI 18K19821 and 22K19833) and by Kyoto University and Toyota Motor Corporation through the joint project Advanced Mathematical Science for Mobility Society.
@@ Appendix A Proofs
Let $\mathcal{H}$ be a non-empty set. A symmetric function $\tilde{k}:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ is a conditionally positive definite kernel on $\mathcal{H}$ if for all $N\in\mathbb{N}$ and $\bm{h}^{(1)},\bm{h}^{(2)},\ldots,\bm{h}^{(N)}\in\mathcal{H}$,
$\sum_{i=1}^{N}\sum_{j=1}^{N}c_{i}c_{j}~{}\tilde{k}\left(\bm{h}^{(i)},\bm{h}^{(j)}\right)\geq 0,$ (21)
with $c_{1},c_{2},\ldots,c_{N}\in\mathbb{R}$ and $\sum_{i=1}^{N}c_{i}=0$.
Theorem 2 (Hilbert space representation of conditionally positive definite kernels (Berg et al., 1984; Schoenberg, 1938; Schölkopf, 2000)).
Let $\mathcal{H}$ be a non-empty set and $\tilde{k}:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ a conditionally positive definite kernel on $\mathcal{H}$ satisfying $\tilde{k}\left(\bm{h},\bm{h}\right)=0$ for all $\bm{h}\in\mathcal{H}$. There exists a Hilbert space $\mathcal{S}$ of real-valued functions on $\mathcal{H}$ and a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{h}^{(1)},\bm{h}^{(1)}\in\mathcal{H}$,
$\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert^{2}=-\tilde{k}\left(\bm{h}^{(1)},\bm{h}^{(2)}\right).$ (22)
@@
Let $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ be a pseudometric. From Assumption 1 and Theorem 2, there exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{h}^{(1)},\bm{h}^{(2)}\in\mathcal{H}$,
$\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert=d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right).$ (23)
@@
$\varepsilon_{2}<\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert<\varepsilon_{1}\iff\varepsilon_{2}<d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)<\varepsilon_{1}.$ (24)
@@
Suppose $\bm{h}^{(0)},\bm{h}^{(1)},\bm{h}^{(2)}\in\mathcal{H}$ and $\tilde{k}:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ is a symmetric function. Then
$k\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)=\dfrac{1}{2}\left[\tilde{k}\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)-\tilde{k}\left(\bm{h}^{(1)},\bm{h}^{(0)}\right)-\tilde{k}\left(\bm{h}^{(0)},\bm{h}^{(2)}\right)+\tilde{k}\left(\bm{h}^{(0)},\bm{h}^{(0)}\right)\right]$ (25)
@@
$D^{2}\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d^{2}(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(1)}\end{subarray}}d^{2}(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(2)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d^{2}(\bm{h},\bm{h^{\prime}})$ (26)
for some pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ and bounded, equinumerous multisets $\bm{H}^{(1)},\bm{H}^{(2)}$. From Assumption 1 and Theorem 3, the pseudometric $d$ has a corresponding positive definite kernel $k:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$. A simple algebraic manipulation and using the fact that $\bm{H}^{(1)},\bm{H}^{(2)}$ are equinumerous results in
$D^{2}\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(1)}\end{subarray}}k(\bm{h},\bm{h^{\prime}})+\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(2)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}k(\bm{h},\bm{h^{\prime}})-2\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}k(\bm{h},\bm{h^{\prime}}).$ (27)
@@
$\left\lVert G\left(\bm{H}^{(1)}\right)-G\left(\bm{H}^{(2)}\right)\right\rVert=D\left(\bm{H}^{(1)},\bm{H}^{(2)}\right),$ (28)
@@
$\varepsilon_{2}<\left\lVert G\left(\bm{H}^{(1)}\right)-G\left(\bm{H}^{(2)}\right)\right\rVert<\varepsilon_{1}\iff\varepsilon_{2}<D\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)<\varepsilon_{1}.$ (30)
Appendix B Experimental set-up
All experiments are conducted on a single NVIDIA® Quadro RTX 6000 (24GB) card using the Deep Graph Library (Wang et al., 2019a) with PyTorch (Paszke et al., 2019) backend. For synthetic datasets, the reported results are obtained from the models at the final epoch across 10 trials with varying seed values. For benchmark datasets, the reported results are obtained from the models with the best validation loss across the 10 trials. The hyperparameters are chosen based on previous results and heuristics without extensive tuning. The codes to reproduce the results may be found at https://github.com/briangodwinlim/SIR-GCN.
B.1 Synthetic datasets
@@
Adopting Brody et al. (2021), the training dataset consists of 4,000 bipartite graphs, each containing $2n$ nodes with randomly assigned attributes and values, while the test dataset comprises 1,000 bipartite graphs with the same configuration. All models utilize a single GNN layer with $4n$ hidden units. A two-layer MLP is also used for GIN and $\sigma$ of SIR-GCN while PNA uses the sum, max, and std aggregators. Model training is performed with the AdamW (Loshchilov & Hutter, 2017) optimizer for over 500 epochs with a batch size of 256 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss.
GraphHeterophily
The training dataset consists of 4,000 directed graphs, each containing a maximum of 50 nodes with randomly assigned labels from one of $c$ classes, while the test dataset comprises 1,000 directed graphs with the same configuration. All models utilize a single GNN layer with $10c$ hidden units and sum pooling as the graph readout function. A feed-forward neural network is also used for GIN while PNA uses the sum, max, and std aggregators. Model training is performed with the AdamW (Loshchilov & Hutter, 2017) optimizer for over 500 epochs with a batch size of 256 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss.
B.2 Benchmark datasets
@@
The datasets are obtained from dgl with data splits (training, validation, test) following Dwivedi et al. (2023). In line with Dwivedi et al. (2023), all models utilize 4 GNN layers with batch normalization and residual connections while constrained with a parameter budget of 100,000. Regularization with weights in $\left\{1\times 10^{-7},1\times 10^{-6},1\times 10^{-5}\right\}$ and dropouts with rates in $\left\{0.1,0.2,0.3\right\}$ are also used to prevent overfitting. The mean, symmetric mean, and max aggregators are used since the sum aggregator is observed to not generalize well to unseen graphs as noted by Veličković et al. (2019). Additionally, sum pooling is used as the graph readout function for ZINC while mean pooling is used for MNIST and CIFAR10. Model training is performed with the AdamW (Loshchilov & Hutter, 2017) optimizer for over a maximum of 500 epochs with a batch size of 128, whenever applicable, and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss. The reported results for other models in Table 3 are obtained from Dwivedi et al. (2023), Corso et al. (2020), and Tailor et al. (2021).
ogbn-arxiv
The dataset is obtained from ogb with data splits (training, validation, test) following Hu et al. (2020). The models utilize a single GNN layer with 256 hidden units, batch normalization, and residual connections. Regularization with weight $1\times 10^{-6}$ and dropouts with rates in increments of 0.1 are also used to prevent overfitting. The symmetric mean aggregator is used along with existing tricks in literature. Model training is performed with the AdamW (Loshchilov & Hutter, 2017) optimizer for over 500 epochs and a learning rate of 0.01 that decays by a factor of 0.5 with patience of 50 epochs based on the training loss. The reported results for other models in Table 4 are obtained from the OGB leaderboard accessible at https://ogb.stanford.edu.
ogbg-molhiv
@@
The dataset is obtained from ogb with data splits (training, validation, test) following Hu et al. (2020). The models utilize a single GNN layer, modified to leverage edge features as described in Appendix D, with 300 hidden units, batch/graph normalization, and residual connections. Regularization with weight $1\times 10^{-7}$ and dropouts with rates in $\left\{0.1,0.4\right\}$ are also used to prevent overfitting. The sum aggregator is used for SIR-GCN aggregation while mean pooling is used as the graph readout function. Model training is performed with the AdamW (Loshchilov & Hutter, 2017) optimizer for over 200 epochs with a batch size of 128 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 20 epochs based on the training loss. The reported results for other models in Table 5 are obtained from the OGB leaderboard accessible at https://ogb.stanford.edu.
Appendix C Model run time

### v4

@@ 6 Conclusion
In summary, the paper provides a new perspective for creating powerful GNNs when the space of node features is uncountable. The key idea is to use pseudometric distances on the space of input to create soft-injective functions such that distinct inputs may produce similar outputs if and only if the distance between the inputs is sufficiently small on some representation. From the theoretical results, SIR-GCN is proposed as a simple and computationally efficient MPNN instance emphasizing contextualized message transformation. Notably, compared to existing MPNN instances, this distinctive feature enables it to learn complex relationships between neighboring nodes and allows it to better handle uncountable node features. Furthermore, the proposed model is also demonstrated to generalize classical GNN methodologies. Despite its simple architectural design and minimal computational requirements, empirical results on synthetic and benchmark datasets underscore the expressivity of SIR-GCN, making it a promising candidate for practical GNN applications. Overall, the paper contributes to GNN literature by theoretically and empirically demonstrating the necessity of both anisotropic and dynamic messages to enhance GNN expressivity. Future works may extend the present framework by considering more complex pseudometric formulations for bounded, equinumerous multisets of $\mathcal{H}$ in Corollary 1. They may also consider a formal analysis of the relationship between contextualized messages and performance on heterophilous graph tasks.
@@ Acknowledgments
This work is supported in part by the Japan Society for the Promotion of Science through the Grants-in-Aid for Scientific Research Program (KAKENHI 18K19821) and in part by Kyoto University and Toyota Motor Corporation through the joint project titled “Advanced Mathematical Science for Mobility Society.”
@@ Appendix A Proofs
Theorem 2 (Hilbert space representation of conditionally positive definite kernels (Schoenberg, 1938; Schölkopf, 2000; Berg et al., 2012)).
Let $\mathcal{H}$ be a non-empty set and $\tilde{k}:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ a conditionally positive definite kernel on $\mathcal{H}$ satisfying $\tilde{k}\left(\boldsymbol{h},\boldsymbol{h}\right)=0$ for all $\boldsymbol{h}\in\mathcal{H}$. There exists a Hilbert space $\mathcal{S}$ of real-valued functions on $\mathcal{H}$ and a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\boldsymbol{h}^{(1)},\boldsymbol{h}^{(1)}\in\mathcal{H}$,
$\left\lVert g\left(\boldsymbol{h}^{(1)}\right)-g\left(\boldsymbol{h}^{(2)}\right)\right\rVert^{2}=-\tilde{k}\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\right).$ (22)
@@
Let $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ be a pseudometric satisfying Assumption 1. By Theorem 2, there exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\in\mathcal{H}$,
$\left\lVert g\left(\boldsymbol{h}^{(1)}\right)-g\left(\boldsymbol{h}^{(2)}\right)\right\rVert=d\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\right).$ (23)
@@
$\varepsilon_{2}<\left\lVert g\left(\boldsymbol{h}^{(1)}\right)-g\left(\boldsymbol{h}^{(2)}\right)\right\rVert<\varepsilon_{1}\iff\varepsilon_{2}<d\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\right)<\varepsilon_{1}.$ (24)
@@
Suppose $\boldsymbol{h}^{(0)},\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\in\mathcal{H}$ and $\tilde{k}:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ is a symmetric function. Then
$k\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\right)\coloneq\dfrac{1}{2}\left[\tilde{k}\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(2)}\right)-\tilde{k}\left(\boldsymbol{h}^{(1)},\boldsymbol{h}^{(0)}\right)-\tilde{k}\left(\boldsymbol{h}^{(0)},\boldsymbol{h}^{(2)}\right)+\tilde{k}\left(\boldsymbol{h}^{(0)},\boldsymbol{h}^{(0)}\right)\right]$ (25)
@@
$D^{2}\left(\boldsymbol{H}^{(1)},\boldsymbol{H}^{(2)}\right)\coloneq\sum_{\begin{subarray}{c}\boldsymbol{h}\in\boldsymbol{H}^{(1)}\\ \boldsymbol{h^{\prime}}\in\boldsymbol{H}^{(2)}\end{subarray}}d^{2}(\boldsymbol{h},\boldsymbol{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\boldsymbol{h}\in\boldsymbol{H}^{(1)}\\ \boldsymbol{h^{\prime}}\in\boldsymbol{H}^{(1)}\end{subarray}}d^{2}(\boldsymbol{h},\boldsymbol{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\boldsymbol{h}\in\boldsymbol{H}^{(2)}\\ \boldsymbol{h^{\prime}}\in\boldsymbol{H}^{(2)}\end{subarray}}d^{2}(\boldsymbol{h},\boldsymbol{h^{\prime}})$ (26)
for some pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ satisfying Assumption 1 and bounded, equinumerous multisets $\boldsymbol{H}^{(1)},\boldsymbol{H}^{(2)}$. By Theorem 3, the pseudometric $d$ has a corresponding positive definite kernel $k:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$. A simple algebraic manipulation and using the fact that $\boldsymbol{H}^{(1)}$ and $\boldsymbol{H}^{(2)}$ are equinumerous results in
$D^{2}\left(\boldsymbol{H}^{(1)},\boldsymbol{H}^{(2)}\right)=\sum_{\begin{subarray}{c}\boldsymbol{h}\in\boldsymbol{H}^{(1)}\\ \boldsymbol{h^{\prime}}\in\boldsymbol{H}^{(1)}\end{subarray}}k(\boldsymbol{h},\boldsymbol{h^{\prime}})+\sum_{\begin{subarray}{c}\boldsymbol{h}\in\boldsymbol{H}^{(2)}\\ \boldsymbol{h^{\prime}}\in\boldsymbol{H}^{(2)}\end{subarray}}k(\boldsymbol{h},\boldsymbol{h^{\prime}})-2\sum_{\begin{subarray}{c}\boldsymbol{h}\in\boldsymbol{H}^{(1)}\\ \boldsymbol{h^{\prime}}\in\boldsymbol{H}^{(2)}\end{subarray}}k(\boldsymbol{h},\boldsymbol{h^{\prime}}).$ (27)
@@
$\left\lVert G\left(\boldsymbol{H}^{(1)}\right)-G\left(\boldsymbol{H}^{(2)}\right)\right\rVert=D\left(\boldsymbol{H}^{(1)},\boldsymbol{H}^{(2)}\right),$ (28)
@@
$\varepsilon_{2}<\left\lVert G\left(\boldsymbol{H}^{(1)}\right)-G\left(\boldsymbol{H}^{(2)}\right)\right\rVert<\varepsilon_{1}\iff\varepsilon_{2}<D\left(\boldsymbol{H}^{(1)},\boldsymbol{H}^{(2)}\right)<\varepsilon_{1}.$ (30)
Appendix B Additional benchmark experiments
Heterophilous Datasets
Heterophilous Datasets (Platonov et al., 2023b) is a collection of node property prediction benchmark datasets for evaluating GNNs under heterophily. In particular, the roman-empire, amazon-ratings, tolokers, and questions datasets have uncountable node features. Meanwhile, the minesweeper dataset has countable node features. The performance metric of roman-empire and amazon-ratings is accuracy while the performance metric of the remaining datasets is ROC-AUC. Platonov et al. (2023b) provides more information regarding the individual datasets.
Table 5 presents the mean and standard deviation of the test performance for SIR-GCN, GCN, GraphSAGE, and GAT across the five benchmark datasets with the experimental set-up closely following that of Platonov et al. (2023b). The test performance for the GraphTransformer (Shi et al., 2021) and heterophilous GNNs—H2GCN (Zhu et al., 2020), CPGNN (Zhu et al., 2021), GPR-GNN (Chien et al., 2021), FSGNN (Maurya et al., 2022), GloGNN (Li et al., 2022), FAGCN (Bo et al., 2021), GBK-GNN (Du et al., 2022), and JacobiConv (Wang & Zhang, 2022)—are also presented as additional baselines. Unsurprisingly, SIR-GCN performs poorly on amazon-ratings, tolokers, and questions as these datasets exhibit near-zero label informativeness (Platonov et al., 2023a) between key node labels and query node labels, which severely limits the ability of SIR-GCN to learn meaningful relationships between neighboring nodes. Despite this inherent challenge, the performance gap in ROC-AUC between SIR-GCN and the top-performing models on tolokers and questions remains minimal, reflecting its robustness. Conversely, SIR-GCN outperforms classical GNNs, GraphTransformer, and all heterophilous GNNs on roman-empire and minesweeper, underscoring its utility even in heterophilous settings. Overall, the results highlight that SIR-GCN remains competitive in heterophilous graph tasks. Future works may further explore its theoretical and empirical properties under heterophily.
Table 5: Test performance on Heterophilous Datasets.
Model roman-empire ($\uparrow$) amazon-ratings ($\uparrow$) minesweeper ($\uparrow$) tolokers ($\uparrow$) questions ($\uparrow$)
GCN 73.69 $\pm$ 0.74 48.70 $\pm$ 0.63 89.75 $\pm$ 0.52 83.64 $\pm$ 0.67 76.09 $\pm$ 1.27
GraphSAGE 85.74 $\pm$ 0.67 53.63 $\pm$ 0.39 93.51 $\pm$ 0.57 82.43 $\pm$ 0.44 76.44 $\pm$ 0.62
GAT 80.87 $\pm$ 0.30 49.09 $\pm$ 0.63 92.01 $\pm$ 0.68 83.70 $\pm$ 0.47 77.43 $\pm$ 1.20
GraphTransformer 86.51 $\pm$ 0.73 51.17 $\pm$ 0.66 91.85 $\pm$ 0.76 83.23 $\pm$ 0.64 77.95 $\pm$ 0.68
H2GCN 60.11 $\pm$ 0.52 36.47 $\pm$ 0.23 89.71 $\pm$ 0.31 73.35 $\pm$ 1.01 63.59 $\pm$ 1.46
CPGNN 63.96 $\pm$ 0.62 39.79 $\pm$ 0.77 52.03 $\pm$ 5.46 73.36 $\pm$ 1.01 65.96 $\pm$ 1.95
GPR-GNN 64.85 $\pm$ 0.27 44.88 $\pm$ 0.34 86.24 $\pm$ 0.61 72.94 $\pm$ 0.97 55.48 $\pm$ 0.91
FSGNN 79.92 $\pm$ 0.56 52.74 $\pm$ 0.83 90.08 $\pm$ 0.70 82.76 $\pm$ 0.61 78.86 $\pm$ 0.92
GloGNN 59.63 $\pm$ 0.69 36.89 $\pm$ 0.14 51.08 $\pm$ 1.23 73.39 $\pm$ 1.17 65.74 $\pm$ 1.19
FAGCN 65.22 $\pm$ 0.56 44.12 $\pm$ 0.30 88.17 $\pm$ 0.73 77.75 $\pm$ 1.05 77.24 $\pm$ 1.26
GBK-GNN 74.57 $\pm$ 0.47 45.98 $\pm$ 0.71 90.85 $\pm$ 0.58 81.01 $\pm$ 0.67 74.47 $\pm$ 0.86
JacobiConv 71.14 $\pm$ 0.42 43.55 $\pm$ 0.48 89.66 $\pm$ 0.40 68.66 $\pm$ 0.65 73.88 $\pm$ 1.16
SIR-GCN 87.67 $\pm$ 0.28 46.73 $\pm$ 0.61 94.12 $\pm$ 0.42 82.85 $\pm$ 0.72 75.33 $\pm$ 1.34
Appendix C Experimental set-up
All experiments are performed on a single NVIDIA® Quadro RTX 6000 (24GB) card using the Deep Graph Library (DGL) (Wang et al., 2019a) with PyTorch (Paszke et al., 2019) backend. For synthetic datasets, the reported results are obtained from the models at the final epoch across 10 trials with varying seed values. For benchmark datasets, the reported results are obtained from the models with the best validation loss across the 10 trials. The codes to reproduce the results may be found at https://github.com/briangodwinlim/SIR-GCN.
C.1 Synthetic datasets
@@
Adopting Brody et al. (2022), the training dataset consists of 4,000 bipartite graphs, each containing $2n$ nodes with randomly assigned attributes and/or values, while the test dataset comprises 1,000 bipartite graphs with the same configuration. All models utilize a single GNN layer with $4n$ hidden units. A two-layer MLP is also used for GIN and $\sigma$ of SIR-GCN while PNA uses the sum, max, and standard deviation aggregators. Model training is performed with the AdamW (Loshchilov & Hutter, 2019) optimizer for 500 epochs with a batch size of 256 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss.
HeteroEdgeCount
The training dataset consists of 4,000 directed graphs, each containing a maximum of 50 nodes with uniformly selected edges using the rand_graph function of DGL and uniformly assigned node labels from one of $c$ classes using the randint function of PyTorch. These measures ensure that the graphs are sufficiently diverse with respect to graph structure and heterophily. Meanwhile, the test dataset comprises 1,000 directed graphs with the same configuration. All models utilize a single GNN layer with $10c$ hidden units and sum pooling as the graph readout function. A feed-forward neural network is also used for GIN while PNA uses the sum, max, and standard deviation aggregators. Model training is performed with the AdamW optimizer for 500 epochs with a batch size of 256 and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss.
C.2 Benchmark datasets
@@
The datasets are obtained from dgl with data splits (training, validation, test) following Dwivedi et al. (2023). In line with Dwivedi et al. (2023), Corso et al. (2020), and Tailor et al. (2022), all models utilize 4 GNN layers with batch normalization and residual connections while constrained to a parameter budget of 100,000. Regularization with weights in $\left\{1\times 10^{-7},1\times 10^{-6},1\times 10^{-5}\right\}$ and dropouts with rates in $\left\{0.1,0.2,0.3\right\}$ are also used to prevent overfitting. The mean, symmetric mean, and max aggregators are used since the sum aggregator is observed to not generalize well to unseen graphs as noted by Veličković et al. (2020). Additionally, sum pooling is used as the graph readout function for ZINC while mean pooling is used for MNIST and CIFAR10. Model training is performed with the AdamW optimizer for a maximum of 500 epochs with a batch size of 128, whenever applicable, and a learning rate of 0.001 that decays by a factor of 0.5 with patience of 10 epochs based on the training loss. The reported results for the other models in Table 3 are obtained from Dwivedi et al. (2023), Corso et al. (2020), and Tailor et al. (2022).
Open Graph Benchmark
@@
The datasets are obtained from ogb, the Open Graph Benchmark Python package, with data splits following Hu et al. (2020a). In line with Corso et al. (2020) and Tailor et al. (2022), the model for ogbn-arxiv and ogbg-molhiv utilizes 3 and 4 GNN layers, respectively, with batch normalization and residual connections while constrained to a parameter budget of 100,000. Weight decays in $\left\{1\times 10^{-4},1\times 10^{-3}\right\}$ and dropouts with a rate of 0.2 are also used to prevent overfitting. The symmetric mean and max aggregators are used with mean pooling as the graph readout function for ogbg-molhiv. Model training is performed with the AdamW optimizer for a maximum of 1000 and 100 epochs, respectively, with a batch size of 64 for ogbg-molhiv and learning rates in $\left\{0.001,0.01\right\}$ that decays by a factor of 0.5 with patience of 40 and 10 epochs, respectively, based on the training loss. The reported results for the other models in Table 4 are obtained from Corso et al. (2020) and Tailor et al. (2022).
Heterophilous Datasets
The datasets are obtained from dgl with data splits following Platonov et al. (2023b). In line with Platonov et al. (2023b), the number of GNN layers is chosen from $\left\{3,5\right\}$, employing residual connections and dropouts with a rate of 0.2. Moreover, the hidden dimension is chosen from $\left\{256,512\right\}$, utilizing batch and layer normalization as well as the symmetric mean and max aggregators. Model training is performed with the AdamW optimizer for a maximum of 1000 epochs and a learning rate of $3\times 10^{-5}$. The reported results for the other models in Table 5 are obtained from Platonov et al. (2023b).
Appendix D Runtime analysis